OpenStudy (anonymous):

Calculate the following integral : \[\int\limits e^\left( ax \right) \cos(bx) dx\]

OpenStudy (mimi_x3):

integration by parts.

OpenStudy (mimi_x3):

u = e^{ax} du / dx = ae^{ax} dv = cos(bx) v = bsin(bx)

OpenStudy (mimi_x3):

\[= e^{ax}{bsin(bx) - \int bsin(bx)*ae^{ax}\]

OpenStudy (mimi_x3):

\[ = e^{ax}{bsin(bx)} - \int bsin(bx)*ae^{ax}\]

OpenStudy (mimi_x3):

integration by parts again.

OpenStudy (mimi_x3):

\[ = e^{ax} bsin(bx) - ab \int sin(bx)e^{ax}\]

OpenStudy (anonymous):

That's what I did, but won't that just go on infinitely?

OpenStudy (mimi_x3):

then you add two integrals together.

OpenStudy (anonymous):

oh, how do you do that?

OpenStudy (mimi_x3):

hmm..what do you get when integrate again? it'll be easier,

OpenStudy (anonymous):

\[(sinbx*e^\left( ax \right))/a - \int\limits (1/a *e^\left( ax \right) * bcosbx) dx\]

OpenStudy (anonymous):

That's just the new integral, so you'd then have to multiply that by the ab and put b*e^(ax)∗sinbx - in front as wel

OpenStudy (mimi_x3):

give me a sec, i did it on paper, i'll upload it.

OpenStudy (anonymous):

ok, great

OpenStudy (mimi_x3):

there's no guarantee, i did it correctly.

OpenStudy (mimi_x3):

and ye zoom in.

OpenStudy (mimi_x3):

control +

OpenStudy (anonymous):

Thank you, I think I understand the method, looks good to me!

OpenStudy (mimi_x3):

:)